# Developed and Quasi-Developed Macro-Scale Heat Transfer in Micro- and Mini-Channels with Arrays of Offset Strip Fins Subject to a Uniform Heat Flux

### v1

@@ 4.1 Onset of the quasi-periodically developed heat transfer regime
Here, the mode’s eigenvalue $\lambda_{T}$ is a constant, while the mode’s amplitude $\Theta$ is spatially periodic along the streamwise coordinate $x_{1}$: $\Theta\left(\bm{x}+\bm{l}_{1}\right)=\Theta\left(\bm{x}\right)$. We therefore define the onset point of quasi-periodically developed heat transfer, $x_{\text{quasi-periodic,T}}$, as the streamwise coordinate $x_{1}$ after which the temperature field agrees with expression (14) within 10% of the temperature difference over a unit cell in the developed region, so $\Theta\exp\left(-\lambda_{T}x_{1}\right)\leq 0.1|\mathrm{\nabla{T}}\bm{\cdot}\bm{l}_{1}|$ for $x_{1}\geq x_{\text{quasi-periodic,T}}$. The former definition of $x_{\text{quasi-periodic,T}}$ is analogous to that of $x_{\text{periodic,T}}$. The extent of the quasi-periodically developed heat transfer region is thus given by $x_{1}\in(x_{\text{quasi-periodic,T}},x_{\text{periodic,T}})$.
@@ 4.2 Eigenvalues and perturbations for quasi-periodically developed heat transfer
From relationship (15) we also learn that the onset point of periodically developed heat transfer will scale as $x_{\text{periodic,T}}\sim 1/\lambda_{T}$ when the peak value of the mode amplitude has no significant effect on the logarithmic factor $\ln\left(\epsilon_{0,T}/\epsilon_{T}\right)$. As we will demonstrate next, the scaling law $x_{\text{periodic,T}}\sim 1/\lambda_{T}$ explains well the influence of the Reynolds number, the Prandtl number and channel aspect ratio on the onset point $x_{\text{periodic,T}}$.
@@ 4.3 Region of quasi-developed macro-scale heat transfer
The macro-scale temperature mode amplitudes $\langle\Theta\rangle_{m}^{f}$ and $\langle\Theta\rangle_{m}^{s}$, in the former expressions (16) for the intrinsic macro-scale temperatures are only a function of the transversal coordinate $x_{2}$ in the channel. In Figures 16 and 16, they are illustrated in their dimensionless form for the same Reynolds number, Prandtl number and geometry as in Figure 14. Their W-like shape strongly resembles the shape of the velocity modes $\langle\bm{U}\rangle_{m}^{f}$ in the quasi-developed flow region because the temperature modes are advected by the same periodic flow field $\bm{u}^{\star}$ as the velocity modes, according to the eigenvalue problem from [55]. This shape reflects that the deviations from the developed macro-scale temperature profile are most pronounced at the center of the channel and at a distance of $l_{1}$ from the side walls.
Theoretically, the magnitude (or peak value) of both amplitudes $\langle\Theta\rangle_{m}^{f}$ and $\langle\Theta\rangle_{m}^{s}$ depends on the specific inlet conditions and channel inlet geometry. Nevertheless, it can be seen from Figures 16 and 16 that the influence of the Reynolds number $Re_{b}$ on the shape of the macro-scale temperature modes is small. As such, their profile can be assumed to be universal for a single channel geometry. For example, for the macro-scale temperature modes in Figures 16 and 16, we can define the following reference profile to characterize their shape at a distance $l_{2}$ away from the side walls:
@@
Closer to the side walls, however, the actual profile diverges from this reference profile due to the imposed no-heat-flux boundary condition and the loss of transversal periodicity in the solid material distribution $\gamma_{s}$. Based on the reference profile (17), the modes for all Reynolds numbers in Figures 16 and 16 can be correlated as $\langle\Theta\rangle_{m}^{f}/\Theta_{\text{ref}}\simeq 0.940Re_{b}^{-1}-7.82e\text{-}5$ and $\langle\Theta\rangle_{m}^{s}/\Theta_{\text{ref}}\simeq 1.15Re_{b}^{-1}-0.00177$, at least within a maximum relative error of 10%. The small coefficients in these correlations verify again our observation that the thermal perturbation size $\epsilon_{0,T}$ is approximately independent of $Re_{b}$, as discussed in Section 4.2.
@@ 4.4 Closure for the quasi-developed heat transfer region
From the asymptotic expression (19), we can determine the section $x_{1}=x_{h}$ after which the actual heat transfer coefficient $h_{b}$ agrees with the developed heat transfer coefficient $h_{\text{unit}}$ within a relative difference $\epsilon_{h}$. However, from our DNS experiments in Section 3.3 and in particular Figure 8, it appears that the exponential correction term on the right hand side of equation (19) adds little accuracy. This is because it decays sufficiently fast, especially when multiplied with the temperature difference $\langle T\rangle_{m}^{f}-\langle T\rangle_{m}^{s}$ to model the macro-scale interfacial heat transfer (5). This is an alternative viewpoint on why we can rely on the developed correlations for $h_{\text{unit}}$ to model the macro-scale interfacial heat transfer $h_{b}$ over nearly the entire channel, except the region near the side walls.
@@
We remark that in principle, the weighting function $m$ of the filter operator $\langle\;\rangle_{m}$ should be matched to both eigenvalues $\lambda_{T}$ and $\lambda$, as discussed in [32, 34]. Otherwise, the exponential terms in (18)-(21) cannot be moved outside the filter operator $\langle\;\rangle_{m}$. Nevertheless, for most cases, the double volume-averaging operator (3) is accurate enough to ensure the validity of the asymptotic forms (18)-(21), especially because we have $\lambda_{T}l_{1}\ll 1$ when the Péclet number is larger than 100 according to Figures 11-13.
@@
with $h_{b}\simeq h_{b,\text{dev}}$. Figure 17 illustrates the validity of this final macro-scale energy balance for a section in the developed region ($x_{1}=18l_{1}$) as well as one in the quasi-developed region ($x_{1}=2l_{1}$). In this figure, the non-dimensional macro-scale velocity, uniform heat flux and macro-scale dispersion are defined such that $\bm{U}_{\text{dev}}^{+}\triangleq\bm{U}_{\text{dev}}/u_{b}$, $q_{b}^{+}\triangleq q_{b}/q_{b}=1$ and $\bm{D}^{+}\triangleq\bm{D}(q_{b}l_{1})/(k_{f}u_{b})$, while $\nabla$ and $\delta_{b}$ denote the dimensionless gradient operator and Diract surface operator based on the reference length $l_{1}$. The main takeaway is that according to the final macro-scale energy balance (23), the developed heat transfer coefficient $h_{b,\text{dev}}$ is the only relevant closure coefficient required to reconstruct the macro-scale temperature profiles of the fluid and solid over almost the entire channel. Although the former heat transfer coefficient is completely known in the core of the channel, where it becomes uniform (i.e. $h_{b,\text{dev}}=h_{b,\text{unit}}$), its spatial behavior closer to the side walls has not yet been investigated. This will be the topic of the next section.
@@ 7 Contributions
The macro-scale description based on the double volume-averaging operation and the validation of its computational framework were carried out by G. Buckinx. G. Buckinx formulated the theoretical framework and modeling equations for the quasi-developed macro-scale heat transfer regime. A. Vangeffelen carried out all numerical heat transfer simulations and post-processing calculations. The results were interpreted by A. Vangeffelen, with input from G. Buckinx regarding the existing literature. The paper was written by A. Vangeffelen and G. Buckinx, with input from C. De Servi, M. R. Vetrano, and M. Baelmans.

### v2

@@ 4.1 Onset of the quasi-periodically developed heat transfer regime
Here, the mode’s eigenvalue $\lambda_{T}$ is a constant, while the mode’s amplitude $\Theta$ is spatially periodic along the streamwise coordinate $x_{1}$: $\Theta\left(\bm{x}+\bm{l}_{1}\right)=\Theta\left(\bm{x}\right)$. We therefore define the onset point of quasi-periodically developed heat transfer, $x_{\text{quasi-periodic,T}}$, as the streamwise coordinate $x_{1}$ after which the temperature field agrees with expression (14) within 10% of the temperature difference over a unit cell in the developed region, so $|T-T_{\text{dev}}-\Theta\exp\left(-\lambda_{T}x_{1}\right)|\leq 0.1|\mathrm{\nabla{T}}\bm{\cdot}\bm{l}_{1}|$ for $x_{1}\geq x_{\text{quasi-periodic,T}}$. The former definition of $x_{\text{quasi-periodic,T}}$ is analogous to that of $x_{\text{periodic,T}}$. The extent of the quasi-periodically developed heat transfer region is thus given by $x_{1}\in(x_{\text{quasi-periodic,T}},x_{\text{periodic,T}})$.
@@ 4.2 Eigenvalues and perturbations for quasi-periodically developed heat transfer
From relationship (15), we also learn that the onset point of periodically developed heat transfer will scale as $x_{\text{periodic,T}}\sim 1/\lambda_{T}$ when the peak value of the mode amplitude has no significant effect on the logarithmic factor $\ln\left(\epsilon_{0,T}/\epsilon_{T}\right)$. As we will demonstrate next, the scaling law $x_{\text{periodic,T}}\sim 1/\lambda_{T}$ explains well the influence of the Reynolds number, the Prandtl number and channel aspect ratio on the onset point $x_{\text{periodic,T}}$.
@@ 4.3 Region of quasi-developed macro-scale heat transfer
The macro-scale temperature mode amplitudes $\langle\Theta\rangle_{m}^{f}$ and $\langle\Theta\rangle_{m}^{s}$, in the former expressions (16) for the intrinsic macro-scale temperatures are only a function of the transversal coordinate $x_{2}$ in the channel. In Figures 17 and 17, they are illustrated in their dimensionless form for the same Reynolds number, Prandtl number and geometry as in Figure 15. Their W-like shape strongly resembles the shape of the velocity modes $\langle\bm{U}\rangle_{m}^{f}$ in the quasi-developed flow region because the temperature modes are advected by the same periodic flow field $\bm{u}^{\star}$ as the velocity modes, according to the eigenvalue problem from [56]. This shape reflects that the deviations from the developed macro-scale temperature profile are most pronounced at the center of the channel and at a distance of $l_{1}$ from the side walls.
Theoretically, the magnitude (or peak value) of both amplitudes $\langle\Theta\rangle_{m}^{f}$ and $\langle\Theta\rangle_{m}^{s}$ depends on the specific inlet conditions and channel inlet geometry. Nevertheless, it can be seen from Figures 17 and 17 that the influence of the Reynolds number $Re_{b}$ on the shape of the macro-scale temperature modes is small. As such, their profile can be assumed to be universal for a single channel geometry. For example, for the macro-scale temperature modes in Figures 17 and 17, we can define the following reference profile to characterize their shape at a distance $l_{2}$ away from the side walls:
@@
Closer to the side walls, however, the actual profile diverges from this reference profile due to the imposed no-heat-flux boundary condition and the loss of transversal periodicity in the solid material distribution $\gamma_{s}$. Based on the reference profile (17), the modes for all Reynolds numbers in Figures 17 and 17 can be correlated as $\langle\Theta\rangle_{m}^{f}/\Theta_{\text{ref}}\simeq 0.940Re_{b}^{-1}-7.82e\text{-}5$ and $\langle\Theta\rangle_{m}^{s}/\Theta_{\text{ref}}\simeq 1.15Re_{b}^{-1}-0.00177$, at least within a maximum relative error of 10%. The small coefficients in these correlations verify again our observation that the thermal perturbation size $\epsilon_{0,T}$ is approximately independent of $Re_{b}$, as discussed in Section 4.2.
@@ 4.4 Closure for the quasi-developed heat transfer region
From the asymptotic expression (19), we can determine the section $x_{1}=x_{h}$ after which the actual heat transfer coefficient $h_{b}$ agrees with the developed heat transfer coefficient $h_{\text{unit}}$ within a relative difference $\epsilon_{h}$. However, from our DNS experiments in Section 3.3 and in particular Figures 8 and 9, it appears that the exponential correction term on the right hand side of equation (19) adds little accuracy. This is because it decays sufficiently fast, especially when multiplied with the temperature difference $\langle T\rangle_{m}^{f}-\langle T\rangle_{m}^{s}$ to model the macro-scale interfacial heat transfer (5). This is an alternative viewpoint on why we can rely on the developed correlations for $h_{\text{unit}}$ to model the macro-scale interfacial heat transfer $h_{b}$ over nearly the entire channel, except the region near the side walls.
@@
We remark that in principle, the weighting function $m$ of the filter operator $\langle\;\rangle_{m}$ should be matched to both eigenvalues $\lambda_{T}$ and $\lambda$, as discussed in [32, 34]. Otherwise, the exponential terms in (18)-(21) cannot be moved outside the filter operator $\langle\;\rangle_{m}$. Nevertheless, for most cases, the double volume-averaging operator (3) is accurate enough to ensure the validity of the asymptotic forms (18)-(21), especially because we have $\lambda_{T}l_{1}\ll 1$ when the Péclet number is larger than 100 according to Figures 12-14.
@@
with $h_{b}\simeq h_{b,\text{dev}}$. Figure 18 illustrates the validity of this final macro-scale energy balance for a section in the developed region ($x_{1}=18l_{1}$) as well as one in the quasi-developed region ($x_{1}=2l_{1}$). In this figure, the non-dimensional macro-scale velocity, uniform heat flux and macro-scale dispersion are defined such that $\bm{U}_{\text{dev}}^{+}\triangleq\bm{U}_{\text{dev}}/u_{b}$, $q_{b}^{+}\triangleq q_{b}/q_{b}=1$ and $\bm{D}^{+}\triangleq\bm{D}(q_{b}l_{1})/(k_{f}u_{b})$, while $\nabla$ and $\delta_{b}$ denote the dimensionless gradient operator and Dirac surface operator based on the reference length $l_{1}$. The main takeaway is that according to the final macro-scale energy balance (23), the developed heat transfer coefficient $h_{b,\text{dev}}$ is the only relevant closure coefficient required to reconstruct the macro-scale temperature profiles of the fluid and solid over almost the entire channel. Although the former heat transfer coefficient is completely known in the core of the channel, where it becomes uniform (i.e. $h_{b,\text{dev}}=h_{b,\text{unit}}$), its spatial behavior closer to the side walls has not yet been investigated. This will be the topic of the next section.
@@ 7 Contributions
The macro-scale description based on the double volume-averaging operation and the validation of its computational framework were carried out by G. Buckinx. G. Buckinx formulated the theoretical framework and modelling equations for the quasi-developed macro-scale heat transfer regime. A. Vangeffelen carried out all numerical heat transfer simulations and post-processing calculations. The results were interpreted by A. Vangeffelen, with input from G. Buckinx regarding the existing literature. The paper was written by A. Vangeffelen and G. Buckinx, with input from C. De Servi, M. R. Vetrano, and M. Baelmans.